The 7th term of an AP is 15 and the fourth is 9. Find the sequence, first term and the common difference

1
Expert's answer
2022-05-30T15:31:09-0400

The common difference could be find according to formula



"d=\frac{a_{m}-a_{n}}{m-n}=\frac{a_{7}-a_{4}}{7-4}=\frac{15-9}{7-4}=\frac{6}{3}=2"

Any term of arithmetic progression is



"a_{n}=a_{1}+(n-1)d"


From here the first term is



"a_{1}=a_{n}-(n-1)d=a_{4}-(4-1)\cdot d=9-(4-1)\cdot 2=3"

Answer: the first term a1=3 and the common difference d=2.